OpenStudy (anonymous):

4. If a person weighs 500 N on Earth, what would she weigh on Jupiter, where the acceleration due to gravity is 26 m/s2? Show all work leading to your answer.

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Any idea? What equations can be used.

OpenStudy (anonymous):

Hint: Your weight changes on the planets, but your mass stays the same.

OpenStudy (anonymous):

yeah M=Fe/Ge

OpenStudy (anonymous):

Yes, F = ma, so solve for the mass, and then use the same formula again for the weight on Jupiter.

OpenStudy (anonymous):

Weight = Force

OpenStudy (anonymous):

so would it look like this 500 N /26 m/s2

OpenStudy (anonymous):

No because 500N is the weight on Earth

OpenStudy (anonymous):

What is the acceleration on earth?

OpenStudy (anonymous):

9.8m/s/s

OpenStudy (anonymous):

Yes, I should specify acceleration due to gravity. 9.8m/s^2

OpenStudy (anonymous):

So you have the force and gravity, now solve for m on earth.

OpenStudy (anonymous):

so 500N i will multiply by 26m/s2

OpenStudy (anonymous):

No, why are you doing that? 500N is your weight on Earth, but you keep using Jupiter's acceleration due to gravity.

OpenStudy (anonymous):

Find the mass, using Fearth = m*gearth

OpenStudy (anonymous):

oh ok Fj=m*g

OpenStudy (anonymous):

Yeah, but you need to find the mass first.

OpenStudy (anonymous):

Mass doesn't change, weight does.

OpenStudy (anonymous):

\[F_{earth} = ma \implies F _{earth}=mg_{earth}\] solve for m.

OpenStudy (anonymous):

F = Mg g = 9.81 m/sec^2 on earth g = 26 m/sec ^2 on Jupiter Woman's weight on Jupiter = 500/9.81 * 26 =1325.18 N

OpenStudy (anonymous):

is that corrrect

OpenStudy (anonymous):

Yeah that's correct, but I hope you understand what is going on in that equation.

OpenStudy (anonymous):

thanks for the help